# All-in-One: Heterogeneous Interaction Modeling for Cold-Start Rating Prediction

### v2

@@ 17       Update model parameters by gradient descent;
UC NeuMF $0.4702_{\small(.0066)}$ $0.7073_{\small(.0034)}$ $0.3713_{\small(.0068)}$ $0.5026_{\small(.0067)}$ $0.6899_{\small(.0045)}$ $0.3947_{\small(.0063)}$ $0.5478_{\small(.0065)}$ $0.6713_{\small(.0052)}$ $0.4391_{\small(.0068)}$
Wide&Deep $0.5189_{\small(.0083)}$ $0.8385_{\small(.0015)}$ $0.4157_{\small(.0102)}$ $0.5561_{\small(.0047)}$ $0.8396_{\small(.0021)}$ $0.4420_{\small(.0058)}$ $0.6248_{\small(.0066)}$ $0.8471_{\small(.0015)}$ $0.5081_{\small(.0093)}$
DeepFM $0.5169_{\small(.0092)}$ $0.8367_{\small(.0079)}$ $0.4123_{\small(.0093)}$ $0.5536_{\small(.0053)}$ $0.8385_{\small(.0073)}$ $0.4384_{\small(.0062)}$ $0.6190_{\small(.0025)}$ $0.8451_{\small(.0063)}$ $0.5019_{\small(.0028)}$
AFN $0.5084_{\small(.0057)}$ $0.8294_{\small(.0032)}$ $0.3998_{\small(.0056)}$ $0.5550_{\small(.0033)}$ $0.8333_{\small(.0029)}$ $0.4340_{\small(.0033)}$ $0.6254_{\small(.0026)}$ $0.8423_{\small(.0018)}$ $0.5028_{\small(.0043)}$
GraphHINGE $0.5180_{\small(.0075)}$ $0.7809_{\small(.0058)}$ $0.4076_{\small(.0071)}$ $0.5819_{\small(.0086)}$ $0.7787_{\small(.0066)}$ $0.4655_{\small(.0086)}$ $0.6462_{\small(.0076)}$ $0.7752_{\small(.0064)}$ $0.5303_{\small(.0094)}$
MetaHIN $0.4392_{\small(.0238)}$ $0.8005_{\small(.0107)}$ $0.3579_{\small(.0245)}$ $0.4892_{\small(.0162)}$ $0.8109_{\small(.0092)}$ $0.3881_{\small(.0179)}$ $0.5644_{\small(.0148)}$ $0.8258_{\small(.0091)}$ $0.4572_{\small(.0191)}$
MAMO $0.4663_{\small(.0053)}$ $0.5905_{\small(.0079)}$ $0.3405_{\small(.0070)}$ $0.5294_{\small(.0038)}$ $0.6119_{\small(.0062)}$ $0.3869_{\small(.0032)}$ $0.6466_{\small(.0063)}$ $0.6531_{\small(.0032)}$ $0.5086_{\small(.0040)}$
TaNP $0.5715_{\small(.0102)}$ $0.8718_{\small(.0041)}$ $0.4728_{\small(.0099)}$ $0.5986_{\small(.0044)}$ $0.8731_{\small(.0026)}$ $0.4899_{\small(.0069)}$ $0.6591_{\small(.0049)}$ $0.8807_{\small(.0020)}$ $0.5495_{\small(.0054)}$
MeLU $0.5093_{\small(.0066)}$ $0.6254_{\small(.0173)}$ $0.4011_{\small(.0149)}$ $0.5787_{\small(.0018)}$ $0.6499_{\small(.0129)}$ $0.4622_{\small(.0107)}$ $0.6882_{\small(.0149)}$ $0.6974_{\small(.0139)}$ $0.5767_{\small(.0167)}$
HIRE $\mathbf{0.6999_{\small(.0184)}}$ $\mathbf{0.9169_{\small(.0145)}}$ $\mathbf{0.6454_{\small(.0309)}}$ $\mathbf{0.7259_{\small(.0042)}}$ $\mathbf{0.9254_{\small(.0176)}}$ $\mathbf{0.6652_{\small(.0144)}}$ $\mathbf{0.7656_{\small(.0246)}}$ $\mathbf{0.9432_{\small(.0223)}}$ $\mathbf{0.7065_{\small(.0211)}}$
IC NeuMF $0.5726_{\small(.0062)}$ $0.7503_{\small(.0045)}$ $0.4982_{\small(.0061)}$ $0.6097_{\small(.0062)}$ $0.7353_{\small(.0048)}$ $0.5356_{\small(.0062)}$ $0.6609_{\small(.0063)}$ $0.7237_{\small(.0055)}$ $0.5933_{\small(.0061)}$
Wide&Deep $0.3006_{\small(.0111)}$ $0.5196_{\small(.0614)}$ $0.1925_{\small(.0174)}$ $0.3751_{\small(.0196)}$ $0.5407_{\small(.0603)}$ $0.2373_{\small(.0328)}$ $0.4681_{\small(.0301)}$ $0.5635_{\small(.0583)}$ $0.3090_{\small(.0533)}$
DeepFM $0.3091_{\small(.0015)}$ $0.5309_{\small(.0649)}$ $0.2012_{\small(.0090)}$ $0.3795_{\small(.0143)}$ $0.5486_{\small(.0609)}$ $0.2451_{\small(.0247)}$ $0.4672_{\small(.0342)}$ $0.5689_{\small(.0537)}$ $0.3134_{\small(.0506)}$
AFN $0.2989_{\small(.0256)}$ $0.4855_{\small(.0878)}$ $0.1891_{\small(.0057)}$ $0.3705_{\small(.0126)}$ $0.5111_{\small(.0856)}$ $0.2308_{\small(.0089)}$ $0.4592_{\small(.0128)}$ $0.5387_{\small(.0780)}$ $0.2965_{\small(.0338)}$
GraphHINGE $0.1428_{\small(.0037)}$ $0.1779_{\small(.0072)}$ $0.0567_{\small(.0017)}$ $0.2087_{\small(.0058)}$ $0.2214_{\small(.0083)}$ $0.0769_{\small(.0027)}$ $0.2971_{\small(.0050)}$ $0.2670_{\small(.0092)}$ $0.1148_{\small(.0019)}$
MetaHIN $0.4369_{\small(.0144)}$ $0.7941_{\small(.0068)}$ $0.3541_{\small(.0159)}$ $0.4934_{\small(.0098)}$ $0.8044_{\small(.0054)}$ $0.3917_{\small(.0122)}$ $0.5624_{\small(.0074)}$ $0.8196_{\small(.0052)}$ $0.4489_{\small(.0106)}$
MAMO $0.4687_{\small(.0132)}$ $0.5942_{\small(.0090)}$ $0.3439_{\small(.0080)}$ $0.5425_{\small(.0023)}$ $0.6194_{\small(.0064)}$ $0.4006_{\small(.0009)}$ $0.6515_{\small(.0040)}$ $0.6590_{\small(.0044)}$ $0.5152_{\small(.0085)}$
TaNP $0.4068_{\small(.0122)}$ $0.7564_{\small(.0060)}$ $0.2720_{\small(.0152)}$ $0.4882_{\small(.0155)}$ $0.7741_{\small(.0065)}$ $0.3339_{\small(.0202)}$ $0.5938_{\small(.0207)}$ $0.7982_{\small(.0077)}$ $0.4383_{\small(.0290)}$
MeLU $0.4893_{\small(.0079)}$ $0.5920_{\small(.0126)}$ $0.3666_{\small(.0172)}$ $0.5708_{\small(.0067)}$ $0.6244_{\small(.0058)}$ $0.4397_{\small(.0035)}$ $0.6860_{\small(.0178)}$ $0.6761_{\small(.0039)}$ $0.5665_{\small(.0139)}$
HIRE $\mathbf{0.5989_{\small(.0161)}}$ $\mathbf{0.8640_{\small(.0162)}}$ $\mathbf{0.5304_{\small(.0307)}}$ $\mathbf{0.6444_{\small(.0116)}}$ $\mathbf{0.8750_{\small(.0188)}}$ $\mathbf{0.5672_{\small(.0186)}}$ $\mathbf{0.7125_{\small(.0284)}}$ $\mathbf{0.8987_{\small(.0222)}}$ $\mathbf{0.6372_{\small(.0260)}}$
U&I C NeuMF $0.5599_{\small(.0061)}$ $0.7059_{\small(.0062)}$ $0.4850_{\small(.0064)}$ $0.6065_{\small(.0053)}$ $0.6936_{\small(.0062)}$ $0.5326_{\small(.0054)}$ $0.6682_{\small(.0036)}$ $0.6885_{\small(.0057)}$ $0.6034_{\small(.0039)}$
Wide&Deep $0.2952_{\small(.0128)}$ $0.5113_{\small(.0610)}$ $0.1857_{\small(.0179)}$ $0.3745_{\small(.0241)}$ $0.5345_{\small(.0586)}$ $0.2337_{\small(.0353)}$ $0.4663_{\small(.0348)}$ $0.5577_{\small(.0559)}$ $0.3039_{\small(.0578)}$
DeepFM $0.3099_{\small(.0073)}$ $0.5286_{\small(.0674)}$ $0.1971_{\small(.0058)}$ $0.3870_{\small(.0074)}$ $0.5467_{\small(.0638)}$ $0.2474_{\small(.0220)}$ $0.4691_{\small(.0330)}$ $0.5651_{\small(.0557)}$ $0.3121_{\small(.0523)}$
AFN $0.2918_{\small(.0290)}$ $0.4749_{\small(.0916)}$ $0.1828_{\small(.0098)}$ $0.3674_{\small(.0183)}$ $0.5016_{\small(.0911)}$ $0.2285_{\small(.0043)}$ $0.4584_{\small(.0116)}$ $0.5317_{\small(.0818)}$ $0.2912_{\small(.0347)}$
GraphHINGE $0.0992_{\small(.0045)}$ $0.1131_{\small(.0062)}$ $0.0335_{\small(.0018)}$ $0.1700_{\small(.0032)}$ $0.1567_{\small(.0055)}$ $0.0552_{\small(.0015)}$ $0.2913_{\small(.0034)}$ $0.2078_{\small(.0043)}$ $0.1248_{\small(.0042)}$
MetaHIN $0.4392_{\small(.0238)}$ $0.8005_{\small(.0107)}$ $0.3579_{\small(.0245)}$ $0.4892_{\small(.0162)}$ $0.8109_{\small(.0092)}$ $0.3881_{\small(.0179)}$ $0.5644_{\small(.0148)}$ $0.8258_{\small(.0091)}$ $0.4572_{\small(.0191)}$
MAMO $0.4789_{\small(.0053)}$ $0.6046_{\small(.0001)}$ $0.3550_{\small(.0020)}$ $0.5391_{\small(.0062)}$ $0.6251_{\small(.0018)}$ $0.3971_{\small(.0026)}$ $0.6480_{\small(.0013)}$ $0.6614_{\small(.0028)}$ $0.5120_{\small(.0069)}$
TaNP $0.4114_{\small(.0112)}$ $0.7663_{\small(.0051)}$ $0.2813_{\small(.0093)}$ $0.4929_{\small(.0100)}$ $0.7820_{\small(.0042)}$ $0.3441_{\small(.0066)}$ $0.5867_{\small(.0074)}$ $0.8016_{\small(.0024)}$ $0.4335_{\small(.0096)}$
MeLU $0.4680_{\small(.0103)}$ $0.5692_{\small(.0060)}$ $0.3393_{\small(.0055)}$ $0.5616_{\small(.0126)}$ $0.6087_{\small(.0072)}$ $0.4210_{\small(.0128)}$ $0.6840_{\small(.0179)}$ $0.6650_{\small(.0094)}$ $0.5556_{\small(.0232)}$
HIRE $\mathbf{0.6030_{\small(.0175)}}$ $\mathbf{0.8693_{\small(.0180)}}$ $\mathbf{0.5362_{\small(.0319)}}$ $\mathbf{0.6627_{\small(.0073)}}$ $\mathbf{0.8837_{\small(.0207)}}$ $\mathbf{0.5904_{\small(.0116)}}$ $\mathbf{0.7419_{\small(.0429)}}$ $\mathbf{0.9140_{\small(.0233)}}$ $\mathbf{0.6764_{\small(.0433)}}$
@@
TABLE IV: Overall Performance in Three Cold-Start Scenarios on $\mathsf{Bookcrossing}$

### v3

@@ 15       Update model parameters by gradient descent;
UC Popularity $58.20_{\small(0.11)}$ $85.40_{\small(0.06)}$ $47.38_{\small(0.16)}$ $64.07_{\small(0.13)}$ $86.40_{\small(0.05)}$ $52.95_{\small(0.17)}$ $72.89_{\small(0.15)}$ $88.00_{\small(0.06)}$ $63.38_{\small(0.20)}$
TransGNN $55.20_{\small(1.80)}$ $56.86_{\small(1.69)}$ $41.29_{\small(2.11)}$ $56.39_{\small(1.80)}$ $57.33_{\small(2.06)}$ $41.40_{\small(2.79)}$ $58.12_{\small(0.48)}$ $57.50_{\small(1.57)}$ $43.40_{\small(1.48)}$
NeuMF $47.02_{\small(0.66)}$ $67.13_{\small(0.52)}$ $37.13_{\small(0.68)}$ $50.26_{\small(0.67)}$ $68.99_{\small(0.45)}$ $39.47_{\small(0.63)}$ $54.78_{\small(0.65)}$ $70.73_{\small(0.34)}$ $43.91_{\small(0.68)}$
Wide&Deep $51.89_{\small(0.83)}$ $83.85_{\small(0.15)}$ $41.57_{\small(1.02)}$ $55.61_{\small(0.47)}$ $83.96_{\small(0.21)}$ $44.20_{\small(0.58)}$ $62.48_{\small(0.66)}$ $84.71_{\small(0.15)}$ $50.81_{\small(0.93)}$
DeepFM $51.69_{\small(0.92)}$ $83.67_{\small(0.79)}$ $41.23_{\small(0.93)}$ $55.36_{\small(0.53)}$ $83.85_{\small(0.73)}$ $43.84_{\small(0.62)}$ $61.90_{\small(0.25)}$ $84.51_{\small(0.63)}$ $50.19_{\small(0.28)}$
AFN $50.84_{\small(0.57)}$ $82.94_{\small(0.32)}$ $39.98_{\small(0.56)}$ $55.50_{\small(0.33)}$ $83.33_{\small(0.29)}$ $43.40_{\small(0.33)}$ $62.54_{\small(0.26)}$ $84.23_{\small(0.18)}$ $50.28_{\small(0.43)}$
GraphHINGE $51.80_{\small(0.75)}$ $77.75_{\small(0.69)}$ $40.76_{\small(0.71)}$ $58.19_{\small(0.86)}$ $78.06_{\small(0.82)}$ $46.55_{\small(0.86)}$ $64.62_{\small(0.76)}$ $78.26_{\small(0.79)}$ $53.03_{\small(0.94)}$
MetaHIN $43.92_{\small(2.38)}$ $80.05_{\small(1.07)}$ $35.79_{\small(2.45)}$ $48.92_{\small(1.62)}$ $81.09_{\small(0.92)}$ $38.81_{\small(1.79)}$ $56.44_{\small(1.48)}$ $82.58_{\small(0.91)}$ $45.72_{\small(1.91)}$
MAMO $46.63_{\small(0.53)}$ $59.05_{\small(0.79)}$ $34.05_{\small(0.70)}$ $52.94_{\small(0.38)}$ $61.19_{\small(0.62)}$ $38.69_{\small(0.32)}$ $64.66_{\small(0.63)}$ $65.31_{\small(0.32)}$ $50.86_{\small(0.40)}$
TaNP $57.15_{\small(1.02)}$ $87.18_{\small(0.41)}$ $47.28_{\small(0.99)}$ $59.86_{\small(0.44)}$ $87.31_{\small(0.26)}$ $48.99_{\small(0.69)}$ $65.91_{\small(0.49)}$ $88.07_{\small(0.20)}$ $54.95_{\small(0.54)}$
MeLU $50.93_{\small(0.66)}$ $62.54_{\small(1.73)}$ $40.11_{\small(1.49)}$ $57.87_{\small(0.18)}$ $64.99_{\small(1.29)}$ $46.22_{\small(1.07)}$ $68.82_{\small(1.49)}$ $69.74_{\small(1.39)}$ $57.67_{\small(1.67)}$
HIRE $\mathbf{69.99_{\small(1.84)}}$ $\mathbf{91.69_{\small(1.45)}}$ $\mathbf{64.54_{\small(3.09)}}$ $\mathbf{72.59_{\small(0.42)}}$ $\mathbf{92.54_{\small(1.76)}}$ $\mathbf{66.52_{\small(1.44)}}$ $\mathbf{76.56_{\small(2.46)}}$ $\mathbf{94.32_{\small(2.23)}}$ $\mathbf{70.65_{\small(2.11)}}$
IC Popularity $57.08_{\small(1.24)}$ $86.14_{\small(0.46)}$ $49.13_{\small(1.15)}$ $60.78_{\small(0.89)}$ $86.26_{\small(0.39)}$ $51.53_{\small(0.99)}$ $67.57_{\small(0.83)}$ $86.94_{\small(0.37)}$ $55.99_{\small(1.03)}$
TransGNN $50.63_{\small(2.36)}$ $50.78_{\small(2.49)}$ $37.96_{\small(2.51)}$ $56.39_{\small(2.01)}$ $53.31_{\small(2.22)}$ $38.36_{\small(2.27)}$ $57.50_{\small(0.83)}$ $54.67_{\small(1.26)}$ $39.40_{\small(1.23)}$
NeuMF $57.26_{\small(0.62)}$ $72.37_{\small(0.55)}$ $49.82_{\small(0.61)}$ $60.97_{\small(0.62)}$ $73.53_{\small(0.48)}$ $53.56_{\small(0.62)}$ $66.09_{\small(0.63)}$ $75.03_{\small(0.45)}$ $59.33_{\small(0.61)}$
Wide&Deep $30.06_{\small(1.11)}$ $51.96_{\small(6.14)}$ $19.25_{\small(1.74)}$ $37.51_{\small(1.96)}$ $54.07_{\small(6.03)}$ $23.73_{\small(3.28)}$ $46.81_{\small(3.01)}$ $56.35_{\small(5.83)}$ $30.90_{\small(5.33)}$
DeepFM $30.91_{\small(0.15)}$ $53.09_{\small(6.49)}$ $20.12_{\small(0.90)}$ $37.95_{\small(1.43)}$ $54.86_{\small(6.09)}$ $24.51_{\small(2.47)}$ $46.72_{\small(3.42)}$ $56.89_{\small(5.37)}$ $31.34_{\small(5.06)}$
AFN $29.89_{\small(2.56)}$ $48.55_{\small(8.78)}$ $18.91_{\small(0.57)}$ $37.05_{\small(1.26)}$ $51.11_{\small(8.56)}$ $23.08_{\small(0.89)}$ $45.92_{\small(1.28)}$ $53.87_{\small(7.80)}$ $29.65_{\small(3.38)}$
GraphHINGE $14.28_{\small(0.37)}$ $17.79_{\small(0.72)}$ $5.67_{\small(0.17)}$ $20.87_{\small(0.58)}$ $22.14_{\small(0.83)}$ $7.69_{\small(0.27)}$ $29.71_{\small(0.50)}$ $26.70_{\small(0.92)}$ $11.48_{\small(0.19)}$
MetaHIN $43.69_{\small(1.44)}$ $79.41_{\small(0.68)}$ $35.41_{\small(1.59)}$ $49.34_{\small(0.98)}$ $80.44_{\small(0.54)}$ $39.17_{\small(1.22)}$ $56.24_{\small(0.74)}$ $81.96_{\small(0.52)}$ $44.89_{\small(1.06)}$
MAMO $46.87_{\small(1.32)}$ $59.42_{\small(0.90)}$ $34.39_{\small(0.80)}$ $54.25_{\small(0.23)}$ $61.94_{\small(0.64)}$ $40.06_{\small(0.09)}$ $65.15_{\small(0.40)}$ $65.90_{\small(0.44)}$ $51.52_{\small(0.85)}$
TaNP $40.68_{\small(1.22)}$ $75.64_{\small(0.60)}$ $27.20_{\small(1.52)}$ $48.82_{\small(1.55)}$ $77.41_{\small(0.65)}$ $33.39_{\small(2.02)}$ $59.38_{\small(2.07)}$ $79.82_{\small(0.77)}$ $43.83_{\small(2.90)}$
MeLU $48.93_{\small(0.79)}$ $59.20_{\small(1.26)}$ $36.66_{\small(1.72)}$ $57.08_{\small(0.67)}$ $62.44_{\small(0.58)}$ $43.97_{\small(0.35)}$ $68.60_{\small(1.78)}$ $67.61_{\small(0.39)}$ $56.65_{\small(1.39)}$
HIRE $\mathbf{59.89_{\small(1.61)}}$ $\mathbf{86.40_{\small(1.62)}}$ $\mathbf{53.04_{\small(3.07)}}$ $\mathbf{64.44_{\small(1.16)}}$ $\mathbf{87.50_{\small(1.88)}}$ $\mathbf{56.72_{\small(1.86)}}$ $\mathbf{71.25_{\small(2.84)}}$ $\mathbf{89.87_{\small(2.22)}}$ $\mathbf{63.72_{\small(2.60)}}$
U&I C Popularity $55.24_{\small(1.71)}$ $86.10_{\small(0.42)}$ $47.13_{\small(1.86)}$ $59.69_{\small(2.29)}$ $86.26_{\small(0.68)}$ $50.54_{\small(2.55)}$ $74.01_{\small(1.70)}$ $86.87_{\small(0.60)}$ $60.16_{\small(2.08)}$
TransGNN $54.38_{\small(1.24)}$ $55.07_{\small(1.45)}$ $42.29_{\small(2.18)}$ $55.21_{\small(1.12)}$ $55.16_{\small(1.23)}$ $42.65_{\small(1.63)}$ $55.63_{\small(0.47)}$ $55.68_{\small(0.62)}$ $46.24_{\small(0.90)}$
NeuMF $55.99_{\small(0.61)}$ $68.85_{\small(0.57)}$ $48.50_{\small(0.64)}$ $60.65_{\small(0.53)}$ $69.36_{\small(0.62)}$ $53.26_{\small(0.54)}$ $66.82_{\small(0.36)}$ $70.59_{\small(0.62)}$ $60.34_{\small(0.39)}$
Wide&Deep $29.52_{\small(1.28)}$ $51.13_{\small(6.10)}$ $18.57_{\small(1.79)}$ $37.45_{\small(2.41)}$ $53.45_{\small(5.86)}$ $23.37_{\small(3.53)}$ $46.63_{\small(3.48)}$ $55.77_{\small(5.59)}$ $30.39_{\small(5.78)}$
DeepFM $30.99_{\small(0.73)}$ $52.86_{\small(6.74)}$ $19.71_{\small(0.58)}$ $38.70_{\small(0.74)}$ $54.67_{\small(6.38)}$ $24.74_{\small(2.20)}$ $46.91_{\small(3.30)}$ $56.51_{\small(5.57)}$ $31.21_{\small(5.23)}$
AFN $29.18_{\small(2.90)}$ $47.49_{\small(9.16)}$ $18.28_{\small(0.98)}$ $36.74_{\small(1.83)}$ $50.16_{\small(9.11)}$ $22.85_{\small(0.43)}$ $45.84_{\small(1.16)}$ $53.17_{\small(8.18)}$ $29.12_{\small(3.47)}$
GraphHINGE $9.92_{\small(0.45)}$ $11.31_{\small(0.62)}$ $3.35_{\small(0.18)}$ $17.00_{\small(0.32)}$ $15.67_{\small(0.55)}$ $5.52_{\small(0.15)}$ $29.13_{\small(0.34)}$ $20.78_{\small(0.43)}$ $12.48_{\small(0.42)}$
MetaHIN $43.92_{\small(2.38)}$ $80.05_{\small(1.07)}$ $35.79_{\small(2.45)}$ $48.92_{\small(1.62)}$ $81.09_{\small(0.92)}$ $38.81_{\small(1.79)}$ $56.44_{\small(1.48)}$ $82.58_{\small(0.91)}$ $45.72_{\small(1.91)}$
MAMO $47.89_{\small(0.53)}$ $60.46_{\small(0.01)}$ $35.50_{\small(0.20)}$ $53.91_{\small(0.62)}$ $62.51_{\small(0.18)}$ $39.71_{\small(0.26)}$ $64.80_{\small(0.13)}$ $66.14_{\small(0.28)}$ $51.20_{\small(0.69)}$
TaNP $41.14_{\small(1.12)}$ $76.63_{\small(0.51)}$ $28.13_{\small(0.93)}$ $49.29_{\small(1.00)}$ $78.20_{\small(0.42)}$ $34.41_{\small(0.66)}$ $58.67_{\small(0.74)}$ $80.16_{\small(0.24)}$ $43.35_{\small(0.96)}$
MeLU $46.80_{\small(1.03)}$ $56.92_{\small(0.60)}$ $33.93_{\small(0.55)}$ $56.16_{\small(1.26)}$ $60.87_{\small(0.72)}$ $42.10_{\small(1.28)}$ $68.40_{\small(1.79)}$ $66.50_{\small(0.94)}$ $55.56_{\small(2.32)}$
HIRE $\mathbf{60.30_{\small(1.75)}}$ $\mathbf{86.93_{\small(1.80)}}$ $\mathbf{53.62_{\small(3.19)}}$ $\mathbf{66.27_{\small(0.73)}}$ $\mathbf{88.37_{\small(2.07)}}$ $\mathbf{59.04_{\small(1.16)}}$ $\mathbf{74.19_{\small(4.29)}}$ $\mathbf{91.40_{\small(2.33)}}$ $\mathbf{67.64_{\small(4.33)}}$
@@
TABLE IV: Overall Performance in 3 Cold-Start Scenarios on $\mathsf{Bookcrossing}$ (%)